OpenStudy (anonymous):

least common denominator 2/3 and 1/5 3/4 and 1/8 5/7, 3/7, and 3/5 1/2, 2/3, and 1/14

OpenStudy (destinymasha):

a least common denominator -- 2 denominators that can go into the same number ex] 2/3 and 4/6 -- 3 can go into 6 , and 6 can go into 6

OpenStudy (destinymasha):

B -- 3/4 and 1/8 -- 4 can go into 8 , and 8 can go into 8

zepdrix (zepdrix):

Here is how I like to do denominators :)\[\Large\bf\sf \frac{3}{4}\quad and\quad \frac{1}{8}\]If we rewrite the 8 as its factors,\[\Large\bf\sf \frac{3}{\color{royalblue}{4}}\quad and\quad \frac{1}{\color{royalblue}{4}\cdot2}\]See how they both have a 4? That means the first fraction needs a factor of 2.\[\Large\bf\sf \frac{2}{2}\cdot\frac{3}{\color{royalblue}{4}}\quad and\quad \frac{1}{\color{royalblue}{4}\cdot2}\]So they can be written as,\[\Large\bf\sf \frac{6}{8}\quad and\quad \frac{1}{8}\]Common denominator of 8. Oh we're not adding them together or anything? :[ Blah my bad, too much detail. Just needed the 8 I guess +_+

OpenStudy (destinymasha):

Thanks for explaining it more :D @zepdrix

OpenStudy (anonymous):

2/3 and 1/5